# This is the extended version of the conference version at FoSSaCS 2022 [BFPZ22] [a] [a] [a]

### v1

@@ 4.2. Checking the global probability-raising condition
for all $t\in\mathbb{R}$. Due to the boundedness of the polyhedron described by conditions (T1)-(T3), (T5), and (T6) and since $z^{\ast}$ lies in the interior of this polyhedron, this means that there must be a value $q\in\mathbb{R}$ such that $(h(z^{\ast}+q\cdot d),g\circ h(z^{\ast}+q\cdot d))$ provides a solution $v$ to equations (T1)-(T3), (T5), and (T6) with an additional 0. By the definition of $g$, this solution furthermore satisfies (T4) and, by equation (6), it still satisfies (T5). This contradicts the choice of the original solution $(x^{\ast},x_{\mathsf{Cause}}^{\ast})$.
@@
We will now suppose that the algorithm guesses the set $X^{\prime}$ as above and show that in this case, it indeed computes a solution. After the variables from $X^{\prime}$ have been removed, $H(y,y_{\mathsf{Cause}})$ can be computed in polynomial time as all the derivatives of $f$ are simple linear expressions in the variables which only require basic arithmetic and can be computed in polynomial time. Likewise, $M$ and $b$ can be computed easily in polynomial time from the original constraints after the guessed variables have been removed. The vectors $c_{0},c_{1},\dots,c_{r}$ describing the solution space to $Mv=b$ can then also be computed in polynomial time.
@@
In summary, the algorithm needs non-determinism to guess the set $X^{\prime}$ of variables which are 0 in a solution to the original constraints with the maximal number of zeroes. All other steps are deterministic polynomial-time computations. Thus, the satisfiability of (S1)-(S5) can be checked in $\mathrm{NP}$.
Memory requirements of schedulers in the original MDP (Proof of Theorem 18). As stated above, every solution to the linear system of inequalities (S1), (S2), and (S3) corresponds to the expected frequencies of state-action pairs of an MR-scheduler in the transformed model satisfying (A1)-(A3). Hence:
@@
Assume (A1)-(A3). Given an MR-scheduler $\mathfrak{U}$ with $\mathrm{Pr}^{\mathfrak{U}}_{\mathcal{M}}(\Diamond\mathsf{Cause})>0$ that violates (GPR), an MR-scheduler $\mathfrak{T}$ with $\mathfrak{T}(s)(\tau)\in\{0,1\}$ for each state $s$ with $\tau\in\mathit{Act}(s)$ that satisfies $\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}}(\Diamond\mathsf{Cause})>0$ and violates (GPR) is computable in polynomial time.
@@
Obviously, $\mathfrak{V}_{y}=\mathfrak{V}$. We now show that at least one of the two MR-schedulers $\mathfrak{V}_{0}$ or $\mathfrak{V}_{1}$ also refutes (GPR). For this, we suppose by contraction that this is not the case, which means that (GPR) holds for both. Let $f:[0,1]\to[0,1]$ be defined by
@@
The scheduler $\mathfrak{T}$ is defined as follows: In all states that do not belong to a MEC $\mathcal{E}$ of $\mathcal{M}$ with $\mathfrak{S}(s_{\mathcal{E}})(\tau)=0$, the behavior of $\mathfrak{T}$ is memoryless: For each state $s$ of $\mathcal{M}$ (and hence of $\mathcal{M}^{\prime}$) that does not belong to a MEC, $\mathfrak{T}(s)=\mathfrak{S}(s)$. For each state $s$ in an end component $\mathcal{E}$ of $\mathcal{M}$ with $\mathfrak{S}(s_{\mathcal{E}})(\tau)=1$, we define $\mathfrak{T}(s)(\tau)=1$. If a MEC $\mathcal{E}$ of $\mathcal{M}$ with $\mathfrak{S}(s_{\mathcal{E}})(\tau)=0$ is entered, $\mathfrak{T}$ makes use of finitely many memory modes as follows: Enumerate the state action pairs $(s,\alpha)$ where $s$ belongs to $\mathcal{E}$, but $\alpha$ does not belong to $\mathcal{E}$, and for which $\mathfrak{S}(s_{\mathcal{E}})(\alpha)>0$ by $(s_{1},\alpha_{1})$, …, $(s_{k},\alpha_{k})$ for some natural number $k$. Further, let $p_{i}\stackrel{{\scriptstyle\text{\rm\tiny def}}}{{=}}\mathfrak{S}(s_{\mathcal{E}})(\alpha_{i})>0$ for all $1\leq i\leq k$. By assumption $\sum_{1\leq i\leq k}p_{i}=1$.
When entering $\mathcal{E}$, the scheduler works in $k$ memory modes $1$, …, $k$ until an action $\alpha$ that does not belong to $\mathcal{E}$ is scheduled starting in memory mode $1$. In each memory mode $i$, $\mathfrak{T}$ follows an MD-scheduler for $\mathcal{E}$ that reaches $s_{i}$ with probability $1$ from all states of $\mathcal{E}$. Once, $s_{i}$ is reached, $\mathfrak{T}$ chooses action $\alpha_{i}$ with probability
@@
Note that this means that $\mathfrak{T}$ leaves $\mathcal{E}$ via $(s_{k},\alpha_{k})$ with probability $1$ if it reaches the last memory mode $k$. As $\mathfrak{T}$ behaves in a memoryless deterministic way in each memory mode, it leaves the end component $\mathcal{E}$ after finitely many steps in expectation. Furthermore, for each $i\leq k$, it leaves $\mathcal{E}$ via $(s_{i},\alpha_{i})$ precisely with probability $(1-\sum_{j<i}p_{j})\cdot q_{i}=p_{i}$. As the behavior of $\mathfrak{S}$ in $\mathit{MEC}(\mathcal{M})$ is hence mimicked by $\mathfrak{T}$ in $\mathcal{M}^{\prime}$, we conclude that the expected frequency of all actions of $\mathcal{M}$ that do not belong to an end component is the same in $\mathcal{M}^{\prime}$ under $\mathfrak{T}$ and in $\mathit{MEC}(\mathcal{M})$ under $\mathfrak{S}$.
As each end component of $\mathcal{M}^{\prime}$ is either left directly via $\tau$ under $\mathfrak{T}$ or after finitely many steps in expectation as just described, the expected frequency of each state-action pair of $\mathcal{M}^{\prime}$ under $\mathfrak{T}$ is finite. In the terminology of [Kal20], the scheduler $\mathfrak{T}$ is transient. By [Kal20, Theorem 9.16], this implies that there is a MR-scheduler $\mathfrak{U}$ for $\mathcal{M}^{\prime}$ under which the expected frequency of all state-action pairs is the same as under $\mathfrak{T}$. So, for this scheduler $\mathfrak{U}$, the expected frequency in $\mathcal{M}^{\prime}$ of all actions $\alpha$ of $\mathcal{M}$ that do not belong to an end component is the same as under $\mathfrak{S}$ in $\mathit{MEC}(\mathcal{M})$.
@@
With these results we can now proof the second main result of this section, Theorem 18.
@@
Solving systems of quadratic inequalities with linear side constraints is NP-hard in general (see, e.g., [GJ79]). For convex problems, in which the associated symmetric matrix occurring in the quadratic inequality has only non-negative eigenvalues, the problem is, however, solvable in polynomial time [KTK80]. Unfortunately, the quadratic constraint system given by (S1)-(S5) is not of this form. We observe that even if $\mathsf{Cause}$ is a singleton $\{c\}$ and the variable $x_{\mathsf{eff}_{\mathsf{unc}}}$ is forced to take a constant value $y$ by (S1)-(S3), i.e., by the structure of the MDP, the inequality (S4) takes the form:
@@ 5.1. Quality measures for causes
Like the MCC, almost all (cf. [Pow11]) of the quality measures studied in the literature are algebraic functions (intuitively speaking, build from polynomials, fractions and root functions) in the entries of the confusion matrix. At the end of this section, we will comment on the computational properties of finding good causes when quality is measured by the infimum over all sensible schedulers of an algebraic function in the entries of the confusion matrix.
@@ 5.2. Computation schemes for the quality measures for fixed cause set
For this section, we assume a fixed PR cause $\mathsf{Cause}$ is given and address the problem to compute its quality values. The first observation is, that all quality measures are preserved by the switch from $\mathcal{M}$ to $\mathcal{M}_{[\mathsf{Cause}]}$ as well as the transformations of $\mathcal{M}_{[\mathsf{Cause}]}$ to an MDP that satisfies conditions (A1)-(A3) of Section 4.2.
@@
$\leqslant$”: Each scheduler for $\mathcal{M}_{[\mathsf{Cause}]}$ can be viewed as a scheduler $\mathfrak{S}$ for $\mathcal{M}$ that behaves as an MD-scheduler minimizing the probability for reaching an effect state from every state in $\mathsf{Cause}$ and we have:
@@
$\geqslant$”: Let $\mathfrak{S}$ be a scheduler of $\mathcal{M}$. Let $\mathfrak{T}=\mathfrak{T}_{\mathfrak{S}}$ be the scheduler of $\mathcal{M}$ that behaves as $\mathfrak{S}$ until the first visit to a state in $\mathsf{Cause}$. As soon as $\mathfrak{T}$ has reached $\mathsf{Cause}$, it behaves as an MD-scheduler minimizing the probability to reach $\mathsf{Eff}$. Recall and coverage under $\mathfrak{T}$ and $\mathfrak{S}$ have the form:
$\begin{array}[]{rclcrcl}\operatorname{\mathit{recall}}_{\mathcal{M}}^{\mathfrak{S}}(\mathsf{Cause})&\ =&\frac{x}{x+q}&&\mathit{covrat}_{\mathcal{M}}^{\mathfrak{S}}(\mathsf{Cause})&\ =&\frac{x}{q}\\ \operatorname{\mathit{recall}}_{\mathcal{M}}^{\mathfrak{T}}(\mathsf{Cause})&=&\frac{y}{y+q}&&\mathit{covrat}_{\mathcal{M}}^{\mathfrak{S}}(\mathsf{Cause})&=&\frac{y}{q}\end{array}$
@@
where $\mathfrak{S}$ ranges over all schedulers with $\mathrm{Pr}_{\mathcal{M}}^{\mathfrak{S}}(\Diamond\mathsf{eff}_{\mathsf{cov}})>0$. Moreover, we have
@@
Obviously, if there is no scheduler $\mathfrak{S}$ as in statement (c) then $\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}}(\Diamond C)>0$. Hence, as a consequence of (ii) we obtain:
@@
where $\mathfrak{S}$ ranges over all schedulers with $\mathrm{Pr}_{\mathcal{M}}^{\mathfrak{S}}(\Diamond V)>0$. To compute these, we rely on a polynomial reduction to the classical stochastic shortest path problem [BT91]. For this, consider the MDP $\mathcal{N}$ arising from $\mathcal{M}$ by adding reset transitions from all terminal states $t\in S\backslash V$ to $\mathsf{init}$. Thus, exactly the $V$-states are terminal in $\mathcal{N}$. $\mathcal{N}$ might contain ECs, which, however, do not intersect with $V$. We equip $\mathcal{N}$ with the weight function that assigns $1$ to all states in $U$ and 0 to all other states. For a scheduler $\mathfrak{T}$ with $\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{N}}(\Diamond V)=1$, let $\mathrm{E}^{\mathfrak{T}}_{\mathcal{N}}(\boxplus V)$ be the expected accumulated weight until reaching $V$ under $\mathfrak{T}$. Let $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)=\inf_{\mathfrak{T}}\mathrm{E}^{\mathfrak{T}}_{\mathcal{N}}(\boxplus V)$ and $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)=\sup_{\mathfrak{T}}\mathrm{E}^{\mathfrak{T}}_{\mathcal{N}}(\boxplus V)$, where $\mathfrak{T}$ ranges over all schedulers with $\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{N}}(\Diamond V)=1$. We can rely on known results [BT91, dA99, BBD+18] to obtain that both $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$ and $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)$ are computable in polynomial time. As $\mathcal{N}$ has only non-negative weights, $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$ is finite and a corresponding MD-scheduler with minimal expectation exists. If $\mathcal{N}$ has an EC containing at least one $U$-state, which is the case iff $\mathcal{M}$ has a scheduler $\mathfrak{S}$ with $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond U)>0$ and $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond V)=0$, then $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)=+\infty$. Otherwise, $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)$ is finite and the maximum is achieved by an MD-scheduler as well.
@@
Consider the MDP $\mathcal{M}$ from Figure 9 with $\mathsf{Eff}=\{\mathsf{eff}_{\mathsf{unc}},\mathsf{eff}_{\mathsf{cov}}\}$ and suppose the task is to compute $\mathit{covrat}(c)=\min/\mathcal{M}{\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}}}$. The construction for the algorithm is depicted in Figure 9 resulting in $\mathcal{N}$, where reset transitions for $\mathsf{noeff}$ and $\mathsf{eff}_{\mathsf{cov}}$ have been added (red edges) and $\mathsf{eff}_{\mathsf{unc}}$ is the only terminal state. The weight function now assigns $1$ to $\mathsf{eff}_{\mathsf{cov}}$ and 0 to all others. By Theorem 33 we have
$\mathit{covrat}(c)=\min/\mathcal{M}{\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}}}=\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus\mathsf{eff}_{\mathsf{unc}}).$
@@
$\mathcal{M}$ has a scheduler $\mathfrak{S}$ with $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond U)>0$ and $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond V)=0$ if and only if the transformed MDP $\mathcal{N}$ in Section 5.2 (Max/min ratios of reachability probabilities for disjoint sets of terminal states) has an EC containing at least one $U$-state. Therefore we then have $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)=+\infty.$ Otherwise, $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)=1/\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V).$
For the following we only consider $\min/\mathcal{M}{U,V}=\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$ since the arguments for the maximum are similar. First we show $\min/\mathcal{M}{U,V}\leqslant\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$. For this, we consider an arbitrary scheduler $\mathfrak{S}$ for $\mathcal{M}$. Let
@@
But this yields $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)\ \geqslant\ \min/\mathcal{M}{U,V}$. we can now rely on known results [BT91, dA99, BBD+18] to compute $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$ and $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)$ in polynomial time.
@@
Applying this framework for $\max/\mathcal{M}{U,V}$ to the f-score we now proof Theorem 31.
@@
We use the simplifying assumptions (A1)-(A3) that can be made due to Lemmata 28 and 29. For $\mathit{fscore}(\mathsf{Cause})$ we have after some straight-forward transformations

### v2

@@ 4.2. Checking the global probability-raising condition
for all $t\in\mathbb{R}$. Due to the boundedness of the polyhedron described by conditions (T1)-(T3), (T5) and (T6) and since $z^{\ast}$ lies in the interior of this polyhedron, this means that there must be a value $q\in\mathbb{R}$ such that $(h(z^{\ast}+q\cdot d),g\circ h(z^{\ast}+q\cdot d))$ provides a solution $v$ to equations (T1)-(T3), (T5) and (T6) with an additional 0. By the definition of $g$, this solution furthermore satisfies (T4) and, by equation (6), it still satisfies (T5). This contradicts the choice of the original solution $(x^{\ast},x_{\mathsf{Cause}}^{\ast})$.
@@
Suppose we guess the set $X^{\prime}$ as above. We show that we then compute a solution. After the variables from $X^{\prime}$ have been removed, $H(y,y_{\mathsf{Cause}})$ can be computed in polynomial time as all the derivatives of $f$ are linear expressions in the variables which require basic arithmetic and can be computed in polynomial time. Likewise, $M$ and $b$ can be computed in polynomial time from the original constraints after the guessed variables have been removed. The vectors $c_{0},c_{1},\dots,c_{r}$ describing the solution space to $Mv=b$ can then also be computed in polynomial time.
@@
In summary, the algorithm needs to guess the set $X^{\prime}$ of variables which are 0 in a solution to the original constraints with the maximal number of zeroes. All other steps are deterministic polynomial-time computations. Thus, satisfiability of (S1)-(S5) can be checked in NP.
Memory requirements of schedulers in the original MDP (Proof of Theorem 18). Every solution to the linear system of inequalities (S1), (S2), and (S3) corresponds to the expected frequencies of state-action pairs of an MR-scheduler in the transformed model satisfying (A1)-(A3). Hence:
@@
Assume (A1)-(A3). Given an MR-scheduler $\mathfrak{U}$ with $\mathrm{Pr}^{\mathfrak{U}}_{\mathcal{M}}(\Diamond\mathsf{Cause})>0$ that violates (GPR), an MR-scheduler $\mathfrak{T}$ with $\mathfrak{T}(s)(\tau)\in\{0,1\}$ for each state $s$ with $\tau\in\mathit{Act}(s)$ that satifies $\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}}(\Diamond\mathsf{Cause})>0$ and violates (GPR) is computable in polynomial time.
@@
Obviously, $\mathfrak{V}_{y}=\mathfrak{V}$. We now show that at least one of the two MR-schedulers $\mathfrak{V}_{0}$ or $\mathfrak{V}_{1}$ also refutes (GPR). For this, we suppose by contraction that this is not the case, which means that (GPR) holds for both. Let $f:[0,1]\to\mathbb{Q}$ be defined by
@@
Let us define the scheduler $\mathfrak{T}$: In all states that do not belong to a MEC $\mathcal{E}$ of $\mathcal{M}$ with $\mathfrak{S}(s_{\mathcal{E}})(\tau)=0$, the behavior of $\mathfrak{T}$ is memoryless: For each state $s$ of $\mathcal{M}$ (and hence of $\mathcal{M}^{\prime}$) that does not belong to a MEC, $\mathfrak{T}(s)=\mathfrak{S}(s)$. For each state $s$ in an end component $\mathcal{E}$ of $\mathcal{M}$ with $\mathfrak{S}(s_{\mathcal{E}})(\tau)=1$, we define $\mathfrak{T}(s)(\tau)=1$. If a MEC $\mathcal{E}$ of $\mathcal{M}$ with $\mathfrak{S}(s_{\mathcal{E}})(\tau)=0$ is entered, $\mathfrak{T}$ makes use of finitely many memory modes as follows: Enumerate the state action pairs $(s,\alpha)$ where $s$ belongs to $\mathcal{E}$, but $\alpha$ does not belong to $\mathcal{E}$, and for which $\mathfrak{S}(s_{\mathcal{E}})(\alpha)>0$ by $(s_{1},\alpha_{1})$, …, $(s_{k},\alpha_{k})$. Further, let $p_{i}\stackrel{{\scriptstyle\text{\rm\tiny def}}}{{=}}\mathfrak{S}(s_{\mathcal{E}})(\alpha_{i})>0$ for all $1\leq i\leq k$. By assumption $\sum_{1\leq i\leq k}p_{i}=1$. When entering $\mathcal{E}$, the scheduler works in $k$ memory modes $1$, …, $k$ until an action $\alpha$ that does not belong to $\mathcal{E}$ is scheduled starting in memory mode $1$. In each memory mode $i$, $\mathfrak{T}$ follows an MD-scheduler for $\mathcal{E}$ that reaches $s_{i}$ with probability $1$ from all states of $\mathcal{E}$. Once, $s_{i}$ is reached, $\mathfrak{T}$ chooses action $\alpha_{i}$ with probability
@@
Now $\mathfrak{T}$ leaves $\mathcal{E}$ via $(s_{k},\alpha_{k})$ with probability $1$ if it reaches the last memory mode $k$. As $\mathfrak{T}$ behaves MD in each mode, it leaves the end component $\mathcal{E}$ after finitely many steps in expectation. Furthermore, for each $i\leq k$, it leaves $\mathcal{E}$ via $(s_{i},\alpha_{i})$ with probability $(1-\sum_{j<i}p_{j})\cdot q_{i}=p_{i}$. Since the behavior of $\mathfrak{S}$ in $\mathit{MEC}(\mathcal{M})$ is mimicked by $\mathfrak{T}$ in $\mathcal{M}^{\prime}$, we conclude that the expected frequency of actions in $\mathcal{M}$ which do not belong to an end component is the same in $\mathcal{M}^{\prime}$ under $\mathfrak{T}$ and in $\mathit{MEC}(\mathcal{M})$ under $\mathfrak{S}$.
The expected frequency of each state-action pair of $\mathcal{M}^{\prime}$ under $\mathfrak{T}$ is finite, since each MEC of $\mathcal{M}^{\prime}$ is left after finitely many steps in expectation. In the terminology of [Kal20], the scheduler $\mathfrak{T}$ is transient. By [Kal20, Theorem 9.16], this implies that there is a MR-scheduler $\mathfrak{U}$ for $\mathcal{M}^{\prime}$ under which the expected frequency of state-action pairs is the same as under $\mathfrak{T}$ and thus the expected frequency in $\mathcal{M}^{\prime}$ of actions $\alpha$ of $\mathcal{M}$ that do not belong to an end component is the same as under $\mathfrak{S}$ in $\mathit{MEC}(\mathcal{M})$.
@@
With these results we can now prove the second main result of this section, Theorem 18, stating that if (GPR) does not hold there is a finite-memory scheduler with two memory cells refuting the GPR condition.
@@
Solving systems of quadratic inequalities with linear side constraints is NP-hard in general (see, e.g., [GJ79]). For convex problems, in which the associated symmetric matrix occurring in the quadratic inequality has only non-negative eigenvalues, the problem is, however, solvable in polynomial time [KTK80]. Unfortunately, the quadratic constraint system describing a scheduler refuting (GPR) given by (S1)-(S5) is not of this form. We observe that even if $\mathsf{Cause}$ is a singleton $\{c\}$ and the variable $x_{\mathsf{eff}_{\mathsf{unc}}}$ is forced to take a constant value $y$ by (S1)-(S3), i.e., by the structure of the MDP, the inequality (S4) takes the form:
@@ 5.1. Quality measures for causes
Like the MCC, almost all (cf. [Pow11]) of the quality measures studied in the literature are algebraic functions (intuitively speaking, built from polynomials, fractions and root functions) in the entries of the confusion matrix. At the end of this section, we will comment on the computational properties of finding good causes when quality is measured by the infimum over all sensible schedulers of an algebraic function in the entries of the confusion matrix.
@@ 5.2. Computation schemes for the quality measures for fixed cause set
For this section, we assume a fixed PR cause $\mathsf{Cause}$ is given and address the problem to compute its quality values. The first observation is, that all quality measures are preserved by the switch from $\mathcal{M}$ to $\mathcal{M}_{[\mathsf{Cause}]}$ as well as the transformations of $\mathcal{M}_{[\mathsf{Cause}]}$ to an MDP that satisfies conditions (A1)-(A3) of Section 4.2. In the following Lemmata 28 and 29 we show that the quality measures $\operatorname{\mathit{recall}},\mathit{covrat}$ and $\mathit{fscore}$ of a fixed $\mathsf{Cause}$ are compatible with the model transformations from section 4. These are, on one hand a transformation to $\mathcal{M}_{[\mathsf{Cause}]}$, which only considers the minimal probability to reach $\mathsf{Eff}$ starting from $\mathsf{Cause}$, and on the other hand a transformation to an MDP $\mathcal{N}$ satisfying (A1)-(A3), which has no end components and has exactly four terminal states $\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}},\mathsf{noeff}_{\mathsf{fp}},\mathsf{noeff}_{\mathsf{tn}}$.
@@
$\leqslant$”: A scheduler for $\mathcal{M}_{[\mathsf{Cause}]}$ can be seen as a scheduler $\mathfrak{S}$ for $\mathcal{M}$ behaving as an MD-scheduler minimizing the reachability probability of $\mathsf{Eff}$ from every state in $\mathsf{Cause}$ and we have:
@@
$\geqslant$”: Let $\mathfrak{S}$ be a scheduler of $\mathcal{M}$. Let $\mathfrak{T}$ be the scheduler of $\mathcal{M}$ that behaves as $\mathfrak{S}$ until the first visit to a state in $\mathsf{Cause}$. As soon as $\mathfrak{T}$ has reached $\mathsf{Cause}$, it behaves as an MD-scheduler minimizing the probability to reach $\mathsf{Eff}$. Recall and coverage under $\mathfrak{T}$ and $\mathfrak{S}$ have the form:
$\begin{array}[]{rclcrcl}\operatorname{\mathit{recall}}_{\mathcal{M}}^{\mathfrak{S}}(\mathsf{Cause})&\ =&\frac{x}{x+q}&&\mathit{covrat}_{\mathcal{M}}^{\mathfrak{S}}(\mathsf{Cause})&\ =&\frac{x}{q}\\ \operatorname{\mathit{recall}}_{\mathcal{M}}^{\mathfrak{T}}(\mathsf{Cause})&=&\frac{y}{y+q}&&\mathit{covrat}_{\mathcal{M}}^{\mathfrak{T}}(\mathsf{Cause})&=&\frac{y}{q}\end{array}$
@@
where $\mathfrak{S}$ ranges over all schedulers with $\mathrm{Pr}_{\mathcal{M}}^{\mathfrak{S}}(\Diamond\mathsf{eff}_{\mathsf{cov}})>0$. Moreover, we can show that we can handle the corner case of $\mathit{fscore}(\mathsf{Cause})=0$.
@@
Obviously, if there is no scheduler $\mathfrak{S}$ as in statement (c) then $\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}}(\Diamond C)>0$. Thus, from (ii) we get:
@@
where $\mathfrak{S}$ ranges over all schedulers with $\mathrm{Pr}_{\mathcal{M}}^{\mathfrak{S}}(\Diamond V)>0$.
To compute these, we rely on a polynomial reduction to the classical stochastic shortest path problem [BT91]. For this, consider the MDP $\mathcal{N}$ arising from $\mathcal{M}$ by adding reset transitions from all terminal states $t\in S\backslash V$ to $\mathsf{init}$. Thus, exactly the $V$-states are terminal in $\mathcal{N}$. $\mathcal{N}$ might contain ECs, which, however, do not intersect with $V$. We equip $\mathcal{N}$ with the weight function that assigns $1$ to all states in $U$ and 0 to all other states. For a scheduler $\mathfrak{T}$ with $\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{N}}(\Diamond V)=1$, let $\mathrm{E}^{\mathfrak{T}}_{\mathcal{N}}(\boxplus V)$ be the expected accumulated weight until reaching $V$ under $\mathfrak{T}$. Let $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)=\inf_{\mathfrak{T}}\mathrm{E}^{\mathfrak{T}}_{\mathcal{N}}(\boxplus V)$ and $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)=\sup_{\mathfrak{T}}\mathrm{E}^{\mathfrak{T}}_{\mathcal{N}}(\boxplus V)$, where $\mathfrak{T}$ ranges over all schedulers with $\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{N}}(\Diamond V)=1$. We can rely on known results [BT91, dA99, BBD+18] to obtain that both $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$ and $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)$ are computable in polynomial time. As $\mathcal{N}$ has only non-negative weights, $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$ is finite and a corresponding MD-scheduler with minimal expectation exists. If $\mathcal{N}$ has an EC containing at least one $U$-state, which is the case iff $\mathcal{M}$ has a scheduler $\mathfrak{S}$ with $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond U)>0$ and $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond V)=0$, then $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)=+\infty$. Otherwise, $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)$ is finite and the maximum is achieved by an MD-scheduler as well.
@@
[…]
@@
$\mathcal{M}$ has a scheduler $\mathfrak{S}$ with $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond U)>0$ and $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond V)=0$ if and only if the transformed MDP $\mathcal{N}$ in Section 5.2 (Max/min ratios of reachability probabilities for disjoint sets of terminal states) has an EC containing at least one $U$-state. Therefore, we then have $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)=+\infty.$ Otherwise, $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)$ is finite.
For the following we only consider $\min/\mathcal{M}{U,V}=\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$ since the arguments for the maximum are similar. First, we show $\min/\mathcal{M}{U,V}\geqslant\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$. For this, we consider an arbitrary scheduler $\mathfrak{S}$ for $\mathcal{M}$. Let
@@
But this yields $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)\ \geqslant\ \min/\mathcal{M}{U,V}$. For $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)\ =\ \max/\mathcal{M}{U,V}$ we use similar arguments. We can now rely on known results [BT91, dA99, BBD+18] to compute $\mathrm{E}^{\min}_{\mathcal{N}}(\boxplus V)$ and $\mathrm{E}^{\max}_{\mathcal{N}}(\boxplus V)$ in polynomial time.
@@
Applying this framework for $\max/\mathcal{M}{U,V}$ to the f-score we now prove Theorem 31.
@@
We use the simplifying assumptions (A1)-(A3) that can be made due to Lemmas 28 and 29. For $\mathit{fscore}(\mathsf{Cause})$ we have after some straight-forward transformations